SSTR= 6750

SSE = 8000

nt = 20Ho: u1=u2=u3=u4

Ha: at least one mean is differnt

At alpha=.05 the null hypothesis

A. was designed incorrectly

B. None of these

C. Should be rejected

D. Should not be rejected
Answer #1

Solution:

Number of treatments = k = 4

d.f.TR = k - 1 = 3

nt = 20

d.f.E = (nt) - k = 20 - 4 = 16

d.f. SS MSS F
Treatment 3 6750 2250 4.5
Error 16 8000 500
Total 79

Using F table , at alpha=.05 , critical value is

F0.05,3,16 = 3.239

4.5 > 3.239

So ,

Reject the null hypothesis

Answer is

C. Should be rejected